OpenStudy (anonymous):

Factor x^3+x^2+2x+5-2x^3-x^2-x-5=0

OpenStudy (lasttccasey):

x^3+x^2+2x+5-2x^3-x^2-x-5=0 The 5's cancel, the x^2's cancel, combine like terms and thus far you have: x-x^3=0 Factor out an "x" for the final answer of: \[x(1-x^2)=0\]

OpenStudy (zehanz):

You can go one step further...\[x(1+x)(1-x)=0\]Now that all this has been done, why not write down the solutions of this equation?